[Cytometry] Canto II issues

Ed Podniesinski flow8775 at roadrunner.com
Sat Sep 13 11:39:45 EDT 2014

Hi Fabio,


I had a similar problem with our CANTO-A (original).

You may have intermittent air bubbles in your Flow Cell.

Our issue was debris stuck in the flow restrictor, .01." p/n 339294,
supplying sheath to the Sheath Plenum. Interesting failure of low sheath
fluid in the Plenum transferring air  to the bubble filter then to the flow

Follow the fluid line from 3-way Valve V9 (on the valve block assembly) to
the Sheath Plenum. Between these two points is the fluid resistor. Sometime
when the instrument was plumbed during manufacture, it is run behind a
mounting bracket (that's what was done on our CANTO-A).

I heard that some guys remove this resistor that controls the slow fill of
the Plenum but I don't.

Do a bubble filter purge and watch if there is air in the purge tubing line
from the bubble filter as you perform this script. 

After a fluidic startup, this line should be solid fluid and no air bubble
in the tubing at all. If there is, suspect a partially plugged fluid
resistor, as they are a cheap part to replace, under $10.

I call it the third sheath filter beside the wet cart .2u sheath filter and
.2u bubble filter.


Did this seem to occur after a normal instrument PM when the sheath filters
are replaced?

I recently have seen the white fibers from the new .2u fluid filters shed
off into the Flow Cell from the bubble filter replacement.

One instrument I was able to wash the contamination out by disassembling the
SIT assembly from the flow cell and washing out with reagent grade water.

The other instrument, one fiber is wedged into the flow cell channel edge
where the quartz pieces mesh to make the flow channel. The core stream is
stable but I'm aware of it being there to monitor during recoupling.

After 23 years I'm finally pre washing any ,2u fluid filter I install for a

Maybe you're supposed to do it anyway (thought you paid for the
cleanliness??)  and I was lucky till now.


The interesting fact is the red laser energy (I'm skipping to say power in
mw)  to the flow cell is the only actual laser line monitored at that point,
entering the flow cell laser beam shaping optics.

The other power/ current values reported are from the lasers themselves
before the beam enters the laser fibers.

Stray 633nm laser light is reflected off one of the prismatic beam expanders
onto a small photo sensor with a red filter. The target on this sensor
assembly is tight and can be sensitive to ambient temp.

As long as your CS&T bead shows a pretty good Levy-Jennings plot of delta
voltage for APC channel then I would say there's no issue with Red laser
power wandering. Typically I see closer to 19mw at the red fiber output.

I think the specification is at least 12mw for the 633nm laser line fiber
output. The bear is to maintain >14 mw out of the 488nm laser fiber over
time. The violet is 25mw or better out of the 405nm fiber. 

That red laser power arbitrary number you see is the result of measured
reflected light, as long as you don't see an issue with the CS&T LJ PMT
voltage plot against APC trend over time, then I would say the service
engineer is right in changing it because of sensor alignment change. 

If it becoming a lower and lower values then see if repositioning the sensor
spikes the values higher to new setting. This is assuming the Red laser
power is not changing as noted by LJ trend.

Also, maybe this sensor just needs a bit of re-centering, as this sensor can
be a nuisance as you are describing.


Ed Podniesinski


Roswell Park Cancer Institute







From: cytometry-bounces at lists.purdue.edu
[mailto:cytometry-bounces at lists.purdue.edu] On Behalf Of Jimenez, Fabio
Sent: Friday, September 12, 2014 12:12 PM
To: Cytometry at lists.purdue.edu
Subject: [Cytometry] Canto II issues


Hi All,


I have a couple of questions regarding a FacsCantoII manufactured in 2013. 


The first one is related to a frequent problem encountered during
acquisition in the FSC-A/SSC-A plot, where the population in FSC-A axis
shifts continually to the left /right to finally settle in the far left. We
have tried to fix the problem by doing the carrousel cleaning with contrad
solution, or flow cell cleaning or long clean but sometimes these strategies
are a hit or miss in resolving this issue. The interesting thing is that if
you take the same sample to another cantoII or other machine (FACSaria) you
don't experience this problem. The BD technician during the scheduled
inspection don't find any issues when checking the machine. Any suggestions
on how to fix this problem or what it is the cause of the problem?


My second question is related to the red laser. From time to time we get the
warning of "red laser power low" and when we call BD support they suggest to
power off and power on the machine to resolve the problem; again it is hit
or a miss solution, usually after sometime the laser gets to the range of
accepted values, but even then you can see that the power fluctuates within
that range and is not as solid as the power for the other lasers (blue and
violet).  The last time that we experienced this problem, the BD technician
in order to fix the problem of low power, reseted the range of acceptable
values (lowered) to match those that red laser was producing. My question
is: is this a valid solution to the problem or is just a way to mask the
problem? What are you recommendations to the problem?


Best regards, 


Fabio Jimenez. 
Research Fellow. 

-------------- next part --------------
HTML attachment scrubbed and removed

More information about the Cytometry mailing list